# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v2

@@ Abstract
LLM inference for popular enterprise use cases, such as summarization, RAG, and code-generation, typically observes orders of magnitude longer prompt lengths than generation lengths. This characteristic leads to high cost of prefill and increased response latency. In this paper, we present SwiftKV, a novel model transformation and distillation procedure specifically designed to reduce the time and cost of processing prompt tokens while preserving high quality of generated tokens. SwiftKV combines three key mechanisms: i) SingleInputKV, which prefills later layers’ KV cache using a much earlier layer’s output, allowing prompt tokens to skip much of the model computation, ii) AcrossKV, which merges the KV caches of neighboring layers to reduce the memory footprint and support larger batch size for higher throughput, and iii) a knowledge-preserving distillation procedure that can adapt existing LLMs for SwiftKV with minimal accuracy impact and low compute and data requirement. For Llama-3.1-8B and 70B, SwiftKV reduces the compute requirement of prefill by 50% and the memory requirement of the KV cache by 62.5% while incurring minimum quality degradation across a wide range of tasks. In the end-to-end inference serving using an optimized vLLM implementation, SwiftKV realizes up to 2×2$\times$2 × higher aggregate throughput and 60% lower time per output token. It can achieve a staggering 560 TFlops/GPU of normalized inference throughput, which translates to 16K tokens/s for Llama-3.1-70B in 16-bit precision on 4$\times$ H100 GPUs. Our training, inference, and model implementations are open-sourced and can be found through https://huggingface.co/collections/Snowflake/swiftkv-models-674f7d7474eb789e185d31cb.
@@ 1 Introduction
Large Language Models (LLMs) are quickly becoming an integral enabler of enterprise applications and offerings, including code and data co-pilots (Chen et al., 2021; Pourreza & Rafiei, 2024), retrieval augmented generation (RAG) (Lewis et al., 2020; Lin et al., 2024), summarization (Pu et al., 2023; Zhang et al., 2024), and agentic workflows (Wang et al., 2024; Schick et al., 2023). While it is clear that LLMs can add value to these applications, the cost and speed of inference determine their practicality. Therefore, improving the aggregate throughput and reducing latency of LLM inference has become an increasingly important topic of interest, with various efforts (Sec. 2) tackling the problem from multiple angles.
@@
In this paper, we take a unique approach to improving LLM inference for enterprise applications based on the key observation that typical enterprise workloads process many more input tokens than output tokens. For example, tasks like code completion, text-to-SQL, summarization and RAG each submit long prompts but produce a small number of generated tokens, and a majority of enterprise LLM use cases in Snowflake incur a 10:1 ratio between prompt and generated tokens.
Based on this observation, we designed SwiftKV, which improves throughput and latency by: i) reducing the computation required to pre-fill the KV cache for input tokens, and ii) enabling memory savings to support larger batch sizes needed to serve LLMs more cost effectively (Sheng et al., 2023; Pope et al., 2022; Yu et al., 2022). SwiftKV (Fig. 1) consists of four key components:
@@
Figure 1: Illustration of SwiftKV with 50% SingleInputKV and 2-way AcrossKV. After distillation, the KV cache of layers 5–8 can all be populated using the hidden state outputs of layer 4. For prefill tokens, the query, attention, and MLP operations of layers 5–8 may be skipped altogether, while decode tokens complete all layers. Existing models may be efficiently adapted for SwiftKV by distilling from the original unmodified model using a small dataset. Model knowledge is preserved by keeping the trainable parameters limited to the Q, K, and V projections of the layers affected by SingleInputKV.
@@
SingleInputKV. SingleInputKV rewires an existing model so that the pre-fill stage during inference can skip a number of later layers in the network, and their KV cache are computed by a single earlier layer. SingleInputKV is motivated by the observation that the output hidden states of the later transformer layers do not change significantly (see Sec. 3.2, also independently discovered by Liu et al. (2024c)). With SingleInputKV, the computation required for pre-fill is reduced by approximately the number of layers skipped. We found that it is possible to skip at least 50% of the layers without significantly impacting the model quality (Sec. 4.2), which translates to a 50% reduction of the pre-fill computation in inference.
AcrossKV. While SingleInputKV reduces the pre-fill computation, it does not reduce the KV cache memory requirement. AcrossKV combines the KV projections from multiple adjacent layers into a single one, and share the KV cache across these layers to reduce its size in memory. AcrossKV allows significant memory savings, which unlocks higher throughput by enabling larger batches during inference. In Sec. 3, we show that AcrossKV can reduce the KV cache size by 25% with less than a 1% quality gap. We also show in ablation studies (Sec. 5) that AcrossKV is compatible with existing KV cache quantization methods, which combine to unlock 62.5% reduction in KV cache size.
Knowledge Recovery. Although SingleInputKV and AcrossKV can be applied to existing LLMs with minimal changes to their architectures, we found that the resulting model parameters should still be adapted to the new architecture to recover their original prediction quality. This can be done via distillation from the original model. With SwiftKV, we found that a lightweight distillation is sufficient, with only a fraction of the model (Q, K, and V projections of the affected layers) trained on 680M tokens, which takes less than 3 hours on a single 8×8$\times$8 × H100 node for Llama-3.1-8B.
SwiftKV Optimized Inference. To realize the computation and memory reductions of SingleInputKV and AcrossKV into end-to-end throughput and latency improvements, we implemented SwiftKV in vLLM (Kwon et al., 2023). Our implementation includes several additional optimizations, including fusing all KV-projections beyond layer $l$ into a single GEMM operation, and integrated memory management needed to lower the KV cache memory footprint achievable via AcrossKV.
@@
SwiftKV increases the aggregate throughput of enterprise workloads by up to 2×2$\times$2 ×, while reducing time-to-first-token (TTFT) and time-per-output-token (TPOT) by up to 50% and 60%, respectively. In fact, for Llama-3.1-70B, SwiftKV can achieve a normalized throughput of 560 TFLops/GPU [^1] Normalized throughput and MFU is based on number of floating point operations in the baseline model.. This is an unprecedented 56.6% MFU utilization for inference (Sec. 4.3). We show that SwiftKV incurs minimal quality degradation averaged across a wide range of tasks (Sec. 4.2), including ARC-Challenge (Clark et al., 2018), Winogrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021).
In addition to these main results, in Sec. 5 we discuss the impact of distillation, datasets, choice of trainable parameters for training SwiftKV. We also present our analysis of the hidden state similarities, and how AcrossKV can be extended and combined with other KV cache compression works. Additionally, we also discuss how SingleInputKV can enable compute savings not just during pre-fill but also during decoding phase.
@@
Lastly, we open-sourced the training and inference code for SwiftKV that can be used to fully reproduce our results, as well as several SwiftKV models that can be used directly by the community at https://huggingface.co/collections/Snowflake/swiftkv-models-674f7d7474eb789e185d31cb.
2 Related Work
Hardware and System Optimizations. Lower-precision quantization like FP8 (Kuzmin et al., 2024) can enable the use of tensor-cores to accelerate inference (Luo et al., 2024). System approaches like PagedAttention (Kwon et al., 2023), Tensor-Parallelism(Shoeybi et al., 2020), Split-Fuse (Holmes et al., 2024; Agrawal et al., 2024), FlashAttention (Dao et al., 2024), and their optimized implementations in TensorRT (NVIDIA, 2019), FasterTransformer (NVIDIA, 2021), vLLM (Kwon et al., 2023), and DeepSpeed-Inference (Aminabadi et al., 2022) enable better parallelization, batching, and scheduling to eliminate performance overheads and achieve better hardware peak utilization without impacting model quality. In contrast, SwiftKV is a model architecture optimization and is complementary to these works.
Sparse attention optimizations. Systems such as ALISA (Zhao et al., 2024) and MInference (Jiang et al., 2024) leverage naturally-occurring sparsity patterns in transformer models to reduce the computation of the quadratic attention operation. Particularly, like SwiftKV, MInference also targets the prefill phase of inference. Sparse attention can be particularly effective for very long sequence lengths (e.g. 100K - 1M tokens) when attention is the dominant operation. In comparison, SwiftKV reduces prefill computation by skipping not just the attention operation, but also the query/output projections and MLP of certain layers. This means that (1) SwiftKV can be more suited for inputs with more moderate lengths when MLP is the dominant operation, and (2) SwiftKV, which either runs attention as-is or skips attention entirely, is complementary to sparse attention methods which are more concerned with the implementation of attention itself.
Memory Compression. A wide range of techniques have been developed to reduce the memory need of inference. Lower-precision quantization techniques like FP8/FP4 can reduce the memory for both KV cache and parameters (Hooper et al., 2024). Attention optimization techniques like MQA (Shazeer, 2019), GQA (Ainslie et al., 2023b), low-rank attention (Chang et al., 2024) also reduce the KV Cache. These approaches are complementary to SwiftKV, which we demonstrate in Sec. 4.2 and Sec. 5.3. Like AcrossKV, MiniCache (Liu et al., 2024b) also considers merging the KV cache of consecutive layers. However, AcrossKV enables consolidating more than just two layers, allowing for higher level of compression, and does not require any token retention strategy where distinct KV caches are stored for special tokens, allowing for simpler implementation.
@@ 2 Related Work
Furthermore, while many of these approaches only focus on reducing the memory, SwiftKV can reduce both the prefill computation (via SingleInputKV) along with memory. As we show in Sec. 5.1, compute reduction rather than memory reduction is crucial for accelerating inference in compute-bound scenarios with sufficient memory, which is common in production with modern GPUs (H100, A100).
@@ 3 SwiftKV: Design and Implementation
Figure 2: LEFT: the input similarity of several smaller scale models (Fig. A.1 in the Appendix shows a similar observation for larger models). RIGHT: The time per forward pass of Llama-3.1-8B. SingleInputKV effectively reduces the forward pass processing time across a range of batch sizes.
@@ 3.1 Preliminaries
In transformer models (Vaswani et al., 2017), attention enables each token to focus on other tokens by comparing queries ($Q$) with keys ($K$) and using values ($V$) to compute the final representation. For a sequence of input tokens $x^{(1)},\dots,x^{(n)}$, the projections are defined as follows: $Q=XW_{Q},\quad K=XW_{K},\quad V=XW_{V}$, where $X\in\mathbb{R}^{n\times d}$ are the input embeddings, and $W_{Q}\in\mathbb{R}^{d\times d_{k}}$ and $W_{K},W_{V}\in\mathbb{R}^{d\times d_{g}}$ are trained model parameters with $d_{g}|d_{k}$. Hereafter, we may also refer to $W_{K}$ and $W_{V}$ as a single matrix $W_{KV}\in\mathbb{R}^{d\times 2d_{k}}$.
During the prefill phase of inference, the model processes the entire input sequence at once, computing $K$ and $V$ for all tokens in parallel (or in chunks in the case of Split-Fuse (Holmes et al., 2024; Agrawal et al., 2024)). This typically occurs when the model handles an initial prompt or context.
During the decoding phase of inference, new tokens are generated one at a time. When predicting the next token, only the query ($Q^{(t+1)}$) for the new token needs to be computed, while the model must attend to the keys and values ($K^{(1)},\dots,K^{(t)}$, $V^{(1)},\dots,V^{(t)}$) of all previously processed tokens.
To optimize efficiency in the decoding phase, KV caching is employed. After processing each token $t$, the newly computed $K^{(t)}$ and $V^{(t)}$ are stored in a cache. For the next token $t+1$, only the new query $Q^{(t+1)}$, key $K^{(t+1)}$, and value $V^{(t+1)}$ are computed. The attention computation will then utilize the cached $K$ and $V$ from all prior tokens, allowing for reduced computational overhead during inference.
3.2 SingleInputKV: Project KV cache from a single layer
Assume the input of $l$-th layer is ${\bf x}_{l}$, and its $i$-th token is ${\bf x}_{l}^{(i)}$. Prior studies (Liu et al., 2024c; Gromov et al., 2024) showed that ${\bf x}_{l}$ becomes more similar as the depth grows. Here, we conduct a similar study.
We compute the average input similarity between $l$-th layer’s input and all remaining layers’ input, i.e.,
@@ 3.2 SingleInputKV: Project KV cache from a single layer
where $L$ is the number of layers in the model and $\mathrm{Similarity({\bf x}_{l},{\bf x}_{j})}$ is the average cosine similarity between all ${\bf x}_{l}^{(i)}$ and ${\bf x}_{j}^{(i)}$ tokens.
We use 50 random training examples from HuggingFaceH4/ultrachat_200k to estimate $\mathrm{SimScore}({\bf x}_{l})$, and the results of Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct are shown in the left of Fig. 2. As the layers get deeper, $\mathrm{SimScore}({\bf x}_{l})$ gets higher. Particularly, around half of the depth, the average similarity of ${\bf x}_{l}$ with ${\bf x}_{>l}$ is above $0.5$ for all models, which shows that the difference of input hidden states are small in deeper layers.
@@
Prefill Compute Reduction. SingleInputKV can enable significant reduction in prefill compute during LLM inference. Originally, all input tokens must be processed by all transformer layers. With SingleInputKV, input tokens [^2] The very last input token still needs to compute all layers to generate the first output token. only need to compute ${\mathbf{W}}_{KV}^{j}{\bf x}_{l+1}$ for layers $j>l$ to generate layer $j$’s KV cache, and all other operations (i.e., QO projections, Attention, and MLP) of layers $j>l$ can be skipped entirely. When prefill computation dominates generated token computation, this reduces the total inference computation to approximately $l/L$. Fig. 1 illustrates the operations skipped by SingleInputKV, and Table 1 shows a more detailed example compute breakdown for Llama-3.1-70B.
@@
Model Vocab K,V Q,O MLP Attn. Total Rel. Baseline 4.3 2.6 22 113 160 302 100% 25% SingleInputKV 4.3 2.6 16 85 120 228 75.5% 50% SingleInputKV 4.3 2.6 11 56 80 154 51.0% 50% SingleInputKV 4.3 1.7 11 56 80 153 50.7% + 4×4$\times$4 × AcrossKV
3.3 AcrossKV: Sharing KV cache for consecutive layers
GQA (Ainslie et al., 2023a), one of the most widely adopted KV cache compression methods, showed that the KV cache can be easily shared within a transformer layer. Later, Liu et al. (2024a) showed that the KV cache can be merged for certain pairs of adjacent layers. AcrossKV extends the ideas to cross-layer KV cache sharing.
Particularly, instead of computing KV cache for all of the remaining layers as shown in equation 2, AcrossKV selectively chooses one layer to compute the KV cache for several consecutive layers and share it within the small group. The key idea is shown in Fig. 1. As AcrossKV can combine multiple layers’ KV caches into a single one rather than just two adjacent layers, it offers higher potential reduction ratio compared to Liu et al. (2024a) while simplifying the implementation to realize the benefits of the KV cache reduction. (See Sec.  2 for more detailed comparison with Liu et al. (2024a)).

### v3

@@ Abstract
LLM inference for enterprise applications, such as summarization, RAG, and code-generation, typically observe much longer prompt than generations, leading to high prefill cost and response latency. We present SwiftKV, a novel model transformation and distillation procedure targeted at reducing the prefill compute (in FLOPs) of prompt tokens while preserving high generation quality. First, SwiftKV prefills later layers’ KV cache using an earlier layer’s output, allowing prompt tokens to skip those later layers. Second, SwiftKV employs a lightweight knowledge-preserving distillation procedure that can adapt existing LLMs with minimal accuracy impact. Third, SwiftKV can naturally incorporate KV cache compression to improve inference performance in low-memory scenarios. Our comprehensive experiments show that SwiftKV can effectively reduce prefill computation by 25–50% across several LLM families while incurring minimum quality degradation. In the end-to-end inference serving, SwiftKV realizes up to 2×2$\times$2 × higher aggregate throughput and 60% lower time per output token. It can achieve a staggering 560 TFlops/GPU of normalized inference throughput, which translates to 16K tokens/s for Llama-3.1-70B. SwiftKV is open-sourced at https://github.com/snowflakedb/arctictraining.
[…]
SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation
Aurick Qiao   Zhewei Yao   Samyam Rajbhandari   Yuxiong He Snowflake AI Research San Mateo, CA, United States Correspondence: aurick.qiao@snowflake.com
@@ 1 Introduction
Large Language Models (LLMs) are now an integral enabler of enterprise applications and offerings, including code and data co-pilots (Chen et al., 2021; Pourreza and Rafiei, 2024), retrieval augmented generation (RAG) (Lewis et al., 2020; Lin et al., 2024), summarization (Pu et al., 2023; Zhang et al., 2024), and agentic workflows (Wang et al., 2024; Schick et al., 2023). However, the cost and speed of inference determine their practicality, and improving the throughput and latency of LLM inference has become increasingly important.
While prior works, such as model pruning Ma et al. (2023); Sreenivas et al. (2024), KV cache compression Hooper et al. (2024); Shazeer (2019); Ainslie et al. (2023b); Chang et al. (2024), and sparse attention Zhao et al. (2024); Jiang et al. (2024), have been developed to accelerate LLM inference, they typically significantly degrade the model quality or work best in niche scenarios, such as low-memory environments or extremely long contexts requests (e.g. >100K tokens). On the other hand, production deployments are often compute-bound rather than memory-bound, and such long-context requests are rare amongst diverse enterprise use cases (e.g. those observed at Snowflake).
@@
In this paper, we take a different approach to improving LLM inference based on the key observation that typical enterprise workloads process more input tokens than output tokens. For example, tasks like code completion, text-to-SQL, summarization, and RAG each submit long prompts but produce fewer output tokens (a 10:1 ratio with average prompt length between 500 and 1000 is observed in our production). In these scenarios, inference throughput and latency are often dominated by the cost of prompt processing (i.e. prefill), and reducing this cost is key to improving their performance.
Based on this observation, we designed SwiftKV, which improves throughput and latency by reducing the prefill computation for prompt tokens. SwiftKV (Fig. 1) consists of three key components:
@@
Figure 1: Illustration of SwiftKV 50% prefill reduction and 2-way AcrossKV. After distillation, the KV cache of layers 5–8 can all be populated using the hidden state outputs of layer 4. For prefill tokens, the query, attention, and MLP operations of layers 5–8 may be skipped, while decode tokens complete all layers. Existing models may be efficiently adapted for SwiftKV by distilling from the original model using a small dataset. Model knowledge is preserved by keeping the trainable parameters limited to the Q, K, and V projections of the layers affected by SwiftKV.
Model transformation.
@@
SwiftKV rewires an existing LLM so that the prefill stage during inference can skip a number of later transformer layers, and their KV cache are computed by the last unskipped layer. This is motivated by the observation that the hidden states of later layers do not change significantly (see Sec. 3.2 and (Liu et al., 2024b)). With SwiftKV, prefill compute is reduced by approximately the number of layers skipped.
Optionally, for low-memory scenarios, we show that the SwiftKV model transformation can naturally incorporate KV cache memory reductions by also merging the KV cache of consecutive skipped layers, which we call AcrossKV.
Knowledge Recovery.
After the SwiftKV transformation is applied to the LLM, its prediction quality is recovered via distillation from the original model. A very lightweight distillation is sufficient, with <10% the model weights (Q, K, and V projections of the skipped layers) trained on <1B tokens, which takes less than 3 hours on 8 H100 GPUs for Llama-3.1-8B-Instruct. In contrast, recent prune-and-distill techniques train the entire pruned model on 10–100B tokens (Tang et al., 2025; Sreenivas et al., 2024).
We show that SwiftKV is effective on diverse architectures, including small models (Llama-3.2-3B-Instruct), large models (Llama-3.1-405B-Instruct), mixture-of-experts and latent attention (Deepseek-V2-Lite-Chat). Remarkably, we found that it is possible for SwiftKV to skip 25–50% of the layers for prompt tokens without significantly impacting the model quality across these scenarios.
Optimized Inference.
@@
To realize SwiftKV into end-to-end throughput and latency improvements, we implemented it in vLLM (Kwon et al., 2023) and SGLang Zheng et al. (2024). SwiftKV increases the throughput of enterprise workloads by up to 2×2$\times$2 ×, while reducing time-to-first-token (TTFT) and time-per-output-token (TPOT) by up to 50% and 60%, respectively. In fact, for Llama-3.1-70B-Instruct, SwiftKV achieves a normalized throughput of 560 TFLOPS/GPU [^1] Normalized throughput and MFU is based on number of floating point operations in the baseline model. at an unprecedented 56.6% MFU utilization for inference (Sec. 4.2). SwiftKV incurs minimal quality degradations (<1–2%) averaged across a wide range of tasks (Sec. 4.1), including ARC-Challenge (Clark et al., 2018), Winogrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021).
@@
We open-sourced the training code for SwiftKV at https://github.com/snowflakedb/arctictraining and and inference code at https://github.com/snowflakedb/arcticinference, as well as several SwiftKV models that can be used directly by the community at https://huggingface.co/collections/Snowflake/swiftkv-models-674f7d7474eb789e185d31cb.
2 Related Works
Model pruning and layer skipping.
Prior works have explored reducing the size and compute footprint of LLMs by pruning their weights, followed by post-training on 10–100B tokens to recover accuracy (Tang et al., 2025; Sreenivas et al., 2024; Xia et al., 2024). Compared to these methods, SwiftKV is focused on reducing prefill compute using a much lighter-weight distillation (<1B tokens). Other works explored adaptively skipping layers without pruning weights, and using little to no post-training Ma et al. (2023); Jaiswal et al. (2024); Men et al. (2024); Yang et al. (2024); Ashkboos et al. (2024). These works reduce compute for prefill and decode tokens alike, and typically can skip up to 25% of the model without significant accuracy degradations. SwiftKV reduces prefill compute, and can skip 25–50% of the model without significant accuracy degradations.
KV cache compression.
@@ 2 Related Works
Quantization techniques like FP8/FP4 can reduce the memory for both KV cache and parameters (Hooper et al., 2024). Attention optimizations like MQA (Shazeer, 2019), GQA (Ainslie et al., 2023b), low-rank attention (Chang et al., 2024) also reduce the KV cache. These approaches are complementary to SwiftKV, which we demonstrate in Sec. 4.1 and Sec. B.1. Furthermore, while many of these approaches only focus on reducing the memory, SwiftKV reduces both the prefill compute and memory (via AcrossKV). As we show in Sec. 5.1, compute reduction is crucial for accelerating LLM inference in compute-bound scenarios with sufficient memory, which is common in production with modern GPUs (e.g., A100, H100).
Sparse attention.
Systems such as ALISA (Zhao et al., 2024) and MInference (Jiang et al., 2024) leverage naturally-occurring sparsity patterns in transformer models to reduce the computation of the quadratic attention operation. Sparse attention can be particularly effective for very long sequence lengths (e.g. 100K–1M tokens) when attention is the dominant operation. In comparison, SwiftKV reduces prefill computation by skipping not just the attention operation, but also the query/output projections and MLP of certain layers. This means that SwiftKV can be more suited for inputs with moderate lengths (e.g. <100K) when MLP is the dominant operation. Additionally, SwiftKV either runs or skips attention operations in their entirety, which makes it orthogonal to existing sparse attention methods.
@@ 3 SwiftKV: Design and Implementation
Figure 2: LEFT: input similarity of smaller models. MID: input similarity of larger models. RIGHT: time per forward pass of Llama-3.1-8B-Instruct. SwiftKV reduces the forward pass processing time across a range of batch sizes.
@@ 3.1 Preliminaries
In transformers (Vaswani et al., 2017), attention enables each token to focus on other tokens by comparing queries ($Q$) with keys ($K$), using values ($V$) to compute the final representation. For a sequence of input tokens $x^{(1)},\dots,x^{(n)}$, the projections are: $Q=XW_{Q}$, $K=XW_{K}$, $V=XW_{V}$, where $X\in\mathbb{R}^{n\times d}$ are the input embeddings, and $W_{Q}\in\mathbb{R}^{d\times d_{k}}$ and $W_{K},W_{V}\in\mathbb{R}^{d\times d_{g}}$ are trained model parameters with $d_{g}|d_{k}$. Hereafter, we may also refer to $W_{K}$ and $W_{V}$ as a single matrix $W_{KV}\in\mathbb{R}^{d\times 2d_{k}}$.
During the prefill phase of inference, the model processes the entire input sequence, computing $K$ and $V$ for all tokens in parallel (or in chunks in the case of Split-Fuse (Holmes et al., 2024; Agrawal et al., 2024)). This typically occurs when the model handles an initial prompt or context.
During the decoding phase of inference, new tokens are generated one at a time. When predicting the next token, only the query ($Q^{(t+1)}$) for the new token needs to be computed, while the model attends to the keys and values ($K^{(1)},\dots,K^{(t)}$, $V^{(1)},\dots,V^{(t)}$) of all previous tokens.
In the decoding phase, KV caching is employed. After processing each token $t$, the newly computed $K^{(t)}$ and $V^{(t)}$ are stored in a cache. For the next token $t+1$, only the new query $Q^{(t+1)}$, key $K^{(t+1)}$, and value $V^{(t+1)}$ are computed. The attention computation will then utilize the cached $K$ and $V$ from all prior tokens, allowing for reduced computational overhead during inference.
3.2 SwiftKV: Project KV cache from one layer
Assume the input of $l$-th layer is ${\bf x}_{l}$, and its $i$-th token is ${\bf x}_{l}^{(i)}$. A key property of LLMs is that ${\bf x}_{l}$ becomes more similar as the depth grows (Liu et al., 2024b; Gromov et al., 2024).
To illustreate, we compute the average input similarity between $l$-th layer’s input and all remaining layers’ input, i.e.,
@@ 3.2 SwiftKV: Project KV cache from one layer
where $L$ is the number of layers and $\mathrm{Similarity({\bf x}_{l},{\bf x}_{j})}$ is the average cosine similarity between all ${\bf x}_{l}^{(i)}$ and ${\bf x}_{j}^{(i)}$.
The results of several models are shown in Fig. 2. Deeper layers have higher $\mathrm{SimScore}({\bf x}_{l})$, and at around half of the depth, the average similarity of ${\bf x}_{l}$ with ${\bf x}_{>l}$ is above $0.5$ for all models, which shows that the difference of input hidden states are small in deeper layers.
@@
Prefill Compute Reduction.
SwiftKV enables significant reduction in prefill compute during LLM inference. Originally, all input tokens must be processed by all transformer layers. With SwiftKV, input tokens [^2] The very last input token still needs to compute all layers to generate the first output token. only need to compute ${\mathbf{W}}_{KV}^{j}{\bf x}_{l+1}$ for layers $j>l$ to generate layer $j$’s KV cache, and all other operations (i.e., QO projections, Attention, and MLP) of layers $j>l$ can be skipped entirely. When prefill computation dominates generated token computation, this reduces the total inference computation to approximately $l/L$. Fig. 1 illustrates the operations skipped by SwiftKV, and Table 1 shows a more detailed example compute breakdown for Llama-3.1-70B-Instruct. We note that decoding tokens still propagate through all layers, so additional decoding heads are not necessary for SwiftKV.
@@
Model Vocab K,V Q,O MLP Attn. Total Rel. Baseline 4.3 2.6 22 113 160 302 100% 25% SwiftKV 4.3 2.6 16 85 120 228 75.5% 50% SwiftKV 4.3 2.6 11 56 80 154 51.0% 50% SwiftKV 4.3 1.7 11 56 80 153 50.7% + 4×4$\times$4 × AcrossKV
3.3 AcrossKV: Share KV cache between layers
GQA (Ainslie et al., 2023a), one of the most widely adopted KV cache compression methods, showed that the KV cache can be shared across attention heads within the same transformer layer. Later, Liu et al. (2024a) showed that the KV cache can be merged for certain pairs of adjacent layers. Although SwiftKV’s main focus is on compute reduction rather than memory reduction, we show that KV cache compression can readily be incorporated with SwiftKV. To do this, SwiftKV is supplemented by AcrossKV, which employs cross-layer KV cache sharing to the skipped layers.
Particularly, instead of computing KV cache for all of the skipped layers as shown in equation 2, AcrossKV selects one layer to compute the KV cache for several consecutive layers and share it within the small group (Fig. 1). AcrossKV can combine more than two layers’ KV caches into a single one, which offers higher potential compression ratios than prior works Liu et al. (2024a) that employ cross-layer KV cache merging, while simplifying its implementation.